Status Report

NASA Space Station On-Orbit Status 31 Mar 2004

By SpaceRef Editor
March 31, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 31 Mar 2004

All ISS systems continue to function nominally, except those noted previously or below. 

After wake-up (1:00am EST), post-sleep activities & breakfast, the crew spent several hours working in the DC-1 docking compartment and Service Module transfer compartment (SM PkhO), checking out the two new Russian Orlan-M spacesuits (#25, #26) for their readiness in the event a contingency EVA should become necessary.  There are also three Orlan EVAs planned for the upcoming Increment 9.  All EVA/suit data files from the checkout were then transferred from the PCS (portable computer system) to SSC (station support computer) for subsequent downlink to the ground.  There were no significant issues.   [After activating the suits, Kaleri and Foale inspected their interior & exterior condition, checked out the suits’ BSS water loop interface units in DC-1 and PkhO, disassembled (scavenged) the Orlan BRTA telemetry units from the old suits (#23, #14) and installed them on the new backpacks, then mounted backup O2 tanks (BK-3) from Lab stowage on the BRTAs.  The Orlans and their BSS interface control units were checked out for proper air/water separation of their cooling system and valve function, measured for proper water flow rate through the LCGs (liquid cooled garments) and tested for leaks of their pressure bladders.  After reconfiguring ISS-internal communications assets for EVA mode, the crew tested the Orlans’ BRTA telemetry system, then its Korona-M/Tranzit-B radio system, using Multimeter (MMTs-01) resistance measurements for continuity and performance checks.  Wearing the medical instrumentation belt (harness) used for the recent EVA-9, Kaleri tested comm and biomedical telemetry via the BSS interface system for vital signs and equipment monitoring.  Afterwards, the DC-1 EVA support panel (POV) was disconnected again from the SM audio/comm subsystem (STTS), and support systems, along with the station audio including Tranzit-B, were reconfigured to their initial conditions.]

Since the 3/29 session of the Russian Uragan earth-imaging program was hampered by too much overcast (>30%), another run was scheduled for Sasha Kaleri today, for focusing the Kodak DCS760 digital camera with 800-mm & 400-mm lenses on new targets of nature and industry environment conditions.  [They included the coastal line of Turkey, coastal area or any region of Adzharia, the Great Caucasus Range, the NE Caspian coastal line, a broad view of the Caspian Sea looking south, and plumes of possible dust storms over the Aral Sea.]

The FE also continued the latest series of Diatomeya ocean research observations and photo & video imagery with the Nikon F5 and Sony DVCAM, today depicting bioproductive waters of the North Atlantic.   [At this time of the year waters in the Gulf Stream delta are distinguished by maximum concentration of phytoplankton.  Green and brownish-green chlorophyll pigment may indicate the presence of structures, which could be of great interest and help in bio-geographical area classification and evaluation of seasonal factors’ impact on bioproductive processes in the North Atlantic.]

CDR/SO Mike Foale performed the periodic testing of the cabin air for combustion products (CO/carbon monoxide & HCl/hydrogen chloride) using three 13P-delivered CSA-CP units (compound specific analyzer/combustion products, #1003, #1009, #1010).

Mike completed the regular routine status checkup of the autonomous Lab payloads.  Sasha attended to the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities).  He also did the weekly inspection of the BRPK-1 water condensate separator, checked out the IP-1 airflow sensors in the various RS hatchways and FGB-to-Node hatchway (last time done: 3/17) and prepared the daily IMS inventory “delta” file.

The CDR completed his 17th weekly filling-out of the Food Frequency Questionnaire (FFQ), which keeps a log of his nutritional intake over time on the medical equipment computer (MEC). 

The Science Officer was thanked by the BCAT (Binary Colloidal Alloy Test) team for yesterday’s excellent teleconference, in particular for relaying his observations on the BCAT phase separating samples.  There is one more photo verification session planned, scheduled for next week.   [Mike also reported that BCAT sample #7, a toxicity level 1 substance, had leaked out of a sample container.  According to the payload developer, the container has two independent seals, tested at 15.3 psi, and is also contained within a Ziploc bag serving as third barrier.]
Foale’s preparation of a new NGL (Next Generation Laptop) yesterday was successful.   [Major PC musical chairs: The ThinkPad A31p, which Mike configured with the new version 5.00 SSC software from the NGSD (Next Generation Support Disk) for the OpsLAN (operations local area network), has now taken over the function of SSC7 as main work station in the U.S. segment (USOS).  The former SSC7 has been moved to replace the SSC2 shell in the SM as new Client at the Central Post, and the former SSC2 (#6075) has become the new Russian LT1 laptop, working with KTsP-1 (central post computer #1).  Mike’s time did not suffice for the optional checkout of the original SSC2’s shell (#6066).]

Planning is underway with the crew for cleaning up onboard stowage, preparatory to the arrival of Expedition 9.  This includes finishing up Progress 13P loading and organizing the equipment stowed behind the ZSRs (zero-G storage racks) in the Lab, scheduled for Friday (4/2).   [Rotation of the selected shelving racks will not have to include the CUP RWS (cupola robotic workstation) as originally thought.]
The Elektron oxygen generator continues to run smoothly in 32 amps mode.  The crew was asked by TsUP/Moscow yesterday to increase the nitrogen (N2) gas pressure in the new BZh liquid unit’s hermetically sealed inside chamber containing the electrolysis unit.  Purpose: to prevent a shutdown of the Elektron if the pressurized capsule reaches the pre-set limit value for low pressure (14.1 psi).   [Using the BPA-M N2 purge assembly, connected to the capsule via a manual valve (VN1), Kaleri increased N2 pressure in the capsule from 14.25 to 17.9 psi, which has since remained unchanged.]

In support of crew health/fitness preparations for the end of the mission, Foale and Kaleri were given the Go to use SLDs (subject load devices) for TVIS treadmill loading during exercise.  When running with a “passive” belt, the SLDs must be worn with the individually set SPDs (subject positioning devices).

The three USOS CMGs (control moment gyroscopes) continue to perform well, with no anomalies and with gimbal rates well within predicted ranges.  The new RS-to-USOS attitude control handover procedure, executed for the first time during the last XPOP-to-LVLH switch on 3/27, worked without hitch.  It is designed to reduce the stress on the CMGs during the handover.  The same procedure will be used next Friday, 4/2, when the station will be slewed back to XPOP (at ~3:40am EST).

Another 3-hr. robotics proficiency activity for the crew with the Canadian SSRMS (Space Station Remote Manipulator System) is planned for tomorrow, including continuation of the external survey/inspection of ISS hardware.   [During inspection of recently acquired photos of the portside forward PVR (photovoltaic radiator)’s base-plate for a known “bubble” more surface irregularities near the bubble and on the starboard side base plate were discovered.  A photo survey of the port base plate near the bubble will be part of tomorrow’s SSRMS activities.]

According to latest word from TsUP/MCC-H ballistics planners, there is a very small possibility that another phasing maneuver may be required for the ISS to fine-tune the 7S/Soyuz TMA-3’s landing location in Northern Kazakhstan.   [The reboost burn, executed with the Progress 13P thrusters, would cost about 40 kg propellants total, including necessary pre- & post-burn attitude maneuvering.]

Planning is underway at TsUP to restore, in the near future, the #2 air conditioner (SKV-2) to operation.   [Due to a misaligned coolant line, the repair of the SKV-2 attempted on 2/13 could not be completed.  To connect the line, some holes will have to be drilled in SM panels, and the current planning deals with the precautions to be taken to avoid sharp edges and keep the resulting drill cuttings from getting into the electronics.]

Today’s CEO (Crew Earth Observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Madras (Chennai), India (nadir pass over this coastal city), Karachi, Pakistan (looking slightly right of track), Muglad Basin Fans, Sudan (general views looking up this basin) were requested to help document surface features), London, England, Great Britain (unusually clear for this time of year.  Weather may hold.  Looking a touch left), Pinacates Biosphere, N Mexico (vegetation patterns on this unique volcanic landscape are of interest.  Looking for the black lava-covered area at nadir, near the mouth of the Colorado River.  Field teams have been doing species counts), and Denver, Colorado (nadir pass).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of today, 2:12pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair not completed; to be tested ASAP — see above note).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 159.4; ppCO2 (mmHg) — 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.7.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 752.79; temperature (deg C) — 22.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 754.82; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 755.02; temperature (deg C) — 23.1; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — 22.1
  • PMA-2:  Shell heater temp (deg C) — 20.0

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking, drag reduction-biased).
  • SM batteries:  Battery #8 is off-line; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #6 is off-line (capacity restoration mode, ROM); battery #3 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3926 kg (8655 lb) as of 3/26/04;  [SM(755) + FGB(2512) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:28am EST [= epoch]):

  • Mean altitude — 365.7 km
  • Apogee — 372.8km
  • Perigee — 358.6 km
  • Period — 91.86 min.
  • Inclination (to Equator) —  51.6295 deg
  • Eccentricity — 0.0010519
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30611

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.